OpenStudy (anonymous):

derivative y=(x+3)^3/x*sqrt(x)

OpenStudy (campbell_st):

rewrite the denominator in index form as x^3/2 u - (x+3)^3 dv/dx= 3(x+3)^2 v = x^3/2 dv/dx = 3/2x^1/2 \[dy/dx = (x^{3/2} \times 3(x+1)^2 - (x + 3)^3 \times 3/2x^{1/2})/(x^{3/2})^2\] I'll let you simplify it

OpenStudy (anonymous):

please help me. i have no idea,how to do it :)

OpenStudy (campbell_st):

just find common factors

OpenStudy (anonymous):

can you explane how to know which function to use in eacht case,to get a right answer?

OpenStudy (campbell_st):

the quotient rule is in the form y = u/v then its a case of dy/xd = (vu' - uv')/v^2

OpenStudy (anonymous):

thanx !